Mason Company projects the following sales for the first three months of the​ year: $ 11,500 in January​; $ 11,600 in February​; and $ 15,800 in March. The company expects 60​% of the sales to be cash and the remainder on account. Sales on account are collected​ 50% in the month of the sale and​ 50% in the following month. The Accounts Receivable account has a zero balance on January 1. Round to the nearest dollar. Read the requirementsLOADING.... Requirement 1. Prepare a schedule of cash receipts for Mason for January​, February​, and March. What is the balance in Accounts Receivable on March 31​? ​(Leave unused and zero balance account cells​ blank, do not enter​ "0".) Cash Receipts from Customers January February March Total Total sales January February March Total Cash Receipts from Customers: Accounts Receivable balance, January 1 January—Cash sales January—Credit sales, collection of January sales in January January—Credit sales, collection of January sales in February February—Cash sales February—Credit sales, collection of February sales in February February—Credit sales, collection of February sales in March March—Cash sales March—Credit sales, collection of March sales in March Total cash receipts from customers Accounts Receivable balance, March 31: March—Credit sales, collection of March sales in April
Answer #1

Answer-1)-

MASON COMPANY
Schedule of Collections from sales
Particulars January February March Total
Projected monthly sales $ (a) 11500 11600 15800 38900
Cash sales (b= 60%*a) 6900 6960 9480 23340
Credit Sales (40%*a) 4600 4640 6320 15560
Cash Collection
50% in month of sale 4600*50%= $2300 $4640*50%= $2320 $6320*50%= $3160 7780
50% in following month of sale 4600*50%= $2300 $4640*50%= $2320 4620
Total cash collections (c) 2300 4620 5480 12400
Total collection (d=b+c) 9200 11580 14960 35740

2)- The balance in Accounts Receivable on March 31 = $6320*50% =$3160.